OpenStudy (anonymous):

I'm stuck on this question in a past paper... 'Simplify (y+5)^6/(y+5)^2'

OpenStudy (anonymous):

Remeber the exponent roles a^m/a^n=a^m-n, so in this case the answer will be: (y+5)^6/(y+5)^2=(y+5)^6-2=(y+5)^4